For the reaction N2O4(g)⇌2NO2(g), the value of K at 25∘C is 7.19×10−3.

Calculate [NO2] at equilibrium when [N2O4]=6.90×10−2mol/L.
Answer #1

N2O4(g)⇌2NO2(g),   the value of K at 25∘C is 7.19×10−3.

[N2O4]=6.90×10−2mol/L.

K   = [NO2]^2/[N2O4]

7.19*10^-3   = [NO2]^2/6.9*10^-2

[NO2]^2   = 7.19*10^-3 *6.9*10^-2

[NO2]^2   = 0.00049611

[NO2]       = 0.0223M >>>>>answer